A. Find the area underneath the given curve.


1.) y= -2x²+8

from x=0 to x=1


2.) y= -x²+7

from x=0 to x=1


B. Find the area enclosed by the given curve, the x-axis , and the given lines.


1.) y=-1/3 x² + 10,

from x=1 and x=3


2.) y= x³-8

from x=-1 to x=2


C. Find the area bounded by the given curve and line.


1.) y= x² + 3 and y=7


2.) y= 2x² and y= 4x + 6


3.) y = -x² + 4x and y= x²


4.) y= x³ - 6x² + 8x and y= x² - 4x



1
Expert's answer
2021-03-19T09:06:58-0400

A. 1) "\int_{0}^{1}{-2x^2+8\ dx}"

"-\frac{2}{3}x^3+8x"

"-\frac{2}{3}+8-0=7\frac{1}{3}"


2) "\int_{0}^{1}{-x^2+1\ dx}"

"-\frac{x^3}{3}+x"

"-\frac{1}{3}+1-0=\frac{2}{3}"


B. 1) "\int_{1}^{3}{-\frac{1}{3}x^2+10\ dx}"

"-\frac{x^3}{9}+10x"

"-\frac{27}{9}+30-\left(-\frac{1}{9}+10\right)=17\frac{1}{9}"


2) "\int_{-1}^{2}{x^3-8\ dx}"

"\frac{x^4}{4}-8x"

"\frac{16}{4}-16-\left(\frac{1}{4}-\ -8\right)=\ -20\frac{1}{4}=20\frac{1}{4}\ square\ units"


C. 1) "y=x^2+3"

"y=7"

"7=x^2+3=>x=\ \pm2"

"\int_{-2}^{2}{7-\left(x^2+3\ \right)dx}"

"4x-\frac{x^3}{3}"

"8-\frac{8}{3}-\left(-8-\ -\frac{8}{3}\right)=10\frac{2}{3}"


2) "y=2x^2"

"y=4x+6"

"2x^2=4x+6"

"x^2-2x-3=0"

"x^2-3x+x-3=0"

"x\left(x-3\right)+\left(x-3\right)=0"

"\left(x+1\right)\left(x-3\right)=0"

"x=\ -1\ or\ x=3"

"\int_{-1}^{3}{4x+6-2x^2\ dx}"

"2x^2+6x-\frac{2}{3}x^3"

"18+18-\frac{54}{3}-\left(2-6--\frac{2}{3}\right)=21\frac{1}{3}"

3) "y=\ -x^2+4x\"

"y=x^2"

"x^2=\ -x^2+4x\"

"2x^2-4x=0"

"x\left(x-2\right)=0"

"x=0\ or\ x=2"

"\int_{0}^{2}{-x^2+4x-x^2\ dx}"

"\int_{0}^{2}{-x^2+4x-x^2\ dx}"

"-\frac{2}{3}x^3+2x^2"

"-\frac{16}{3}+8=2\frac{2}{3}"


4) "y=x^3-6x^2"

"y=x^2-4x"

"x^3-7x^2+4x=0"

"x\left(x^2-7x+4\right)=0=>x=0"

"x^2-7x+4=0"

"\frac{7\ \pm\ \sqrt{49-16}}{2}=>x=\ 0.628\ or\ 6.372"

"x\ intercepts"

"{for\ y=x^2-4x\ ,\ \ x}^2-4x=0=>x=0\ or\ x=4"

"for\ y=\ x^3-6x^2, \ x^3-6x^2=0=>x=0\ or\ x=6"

"|\int_{0}^{6}{x^3-6x^2}dx|-|\int_{0}^{4}{x^2-4x\ dx\ |}"

"+\ \int_{4}^{6.372}{x^2-4x\ dx\ }-\ \int_{6}^{6.372}{x^3-6x^2\ dx\ }"

"\left|\frac{x^4}{4}-2x^3\right|0,6-\left|\frac{x^3}{3}-2x^2\right|0,\ 4"

"+\left(\frac{x^3}{3}-2x^2\right),\ 4,\ 6.372-\left(\frac{x^4}{4}-2x^3\right),\ 6,\ 6.372"

"108-10.667+15.701-\ 2.702=110.332"